# 3DXTalker: Unifying Identity, Lip Sync, Emotion, and Spatial Dynamics in Expressive 3D Talking Avatars

### v2

@@ 3.2 Unified Audio-rich Framework
Identity Head. This branch is responsible for predicting the shape $\{\boldsymbol{\beta}\}$ and detail $\{\boldsymbol{\delta}\}$ parameters that define the identity of the generated talking head. With the global latent representation $\mathbf{H}_{t}$ extracted, the identity head employs lightweight self-attention layers $\mathcal{D}_{id}$ to further form identity-related features and subsequently passes them through an MLP layer to produce the velocity field of shape and detail ($\hat{\boldsymbol{v}}^{\beta}_{t}\in\mathbb{R}^{N\times 100}$ and $\hat{\boldsymbol{v}}^{\delta}\in\mathbb{R}^{N\times 128}$), defined as:
@@
Pose Head with Amplitude Embeddings. To achieve precise and responsive control of jaw motion and head rotation, we extract frame-wise amplitude features $\mathbf{A}_{amp}$ from the driving audio. Specifically, we first compute the amplitude envelope of the waveform and then apply frame-level window averaging to obtain a temporally aligned amplitude sequence that reflects speech intensity and rhythmic variations. These amplitude cues are injected into the pose branch via a cross-attention module $\mathcal{D}_{pose}$, where $\mathbf{H}_{t}$ provides motion context and $\mathbf{A}_{amp}$ provides audio-driven modulation. The resulting features are decoded by an MLP to produce the velocity fields for jaw pose $\hat{\boldsymbol{v}}^{\theta^{j}}_{t}\in\mathbb{R}^{N\times 3}$ and global head rotation $\hat{\boldsymbol{v}}^{\theta^{g}}_{t}\in\mathbb{R}^{N\times 3}$:
@@
Expression Head with Emotion Embeddings. To enable emotionally coherent facial expressions, we extract frame-wise emotion embeddings $\mathbf{A}_{emo}$ from the driving audio using emotion2vec [30]. These embeddings capture subtle affective cues (e.g., happiness, sadness, anger) embedded in speech and are temporally aligned with the latent representation. The expression head injects $\mathbf{A}_{emo}$ into the motion representation via a cross-attention module $\mathcal{D}_{exp}$, followed by a MLP that predicts the expression velocity field $\hat{\boldsymbol{v}}^{\psi}_{t}\in\mathbb{R}^{N\times 50}$, represented as:
@@
Flow-matching Inference. During inference, the velocity fields from all heads are concatenated ($\boldsymbol{\hat{v}}_{t}=\text{Concat}(\,\hat{\boldsymbol{v}}^{\beta}_{t},\;\hat{\boldsymbol{v}}^{\psi}_{t},\;\hat{\boldsymbol{v}}^{\theta}_{t},\;\hat{\boldsymbol{v}}^{\delta}_{t}\,)$) to form a unified displacement field $\boldsymbol{\varepsilon}_{t}$, which is iteratively updated to $T_{\text{inf}}$ steps. The final FLAME parameters are obtained by adding this displacement to the reference:
@@
Flow-matching Training. During training, we model the continuous evolution of FLAME parameters as a straight flow between the initial latent $\boldsymbol{\varepsilon}_{0}$ and the target displacement $\mathbf{X}_{\Delta}=\mathbf{X}-\mathbf{X}_{ref}$. Given a randomly sampled $t\sim\mathcal{U}(0,1)$, the target velocity is defined as the linear interpolant, and the flow matching objective supervises the model-predicted velocity $\hat{\boldsymbol{v}}(t)$ as follows:
This formulation directly aligns the model with the continuous flow that transports the latent state to the parameter space, enabling stable and temporally smooth synthesis. These designs allow 3DXTalker to achieve identity consistency, accurate lip synchronization, rich emotional expressions, and natural head motion within a unified framework.
@@
Table 1: Quantitative evaluation with seven baseline models over 200 videos. “ID Gen." denotes whether the model supports identity-referred generalization. Our method achieves better performance on 3D-related metrics, demonstrating improved geometric accuracy and temporal coherence, while it performs well on most 2D perceptual metrics. The BA score assesses the rhythmic alignment of head motions with the driving audio. The MR is computed based on voting results from 74 voting responses. Throughput was measured on an NVIDIA RTX 5090 GPU. The best results are highlighted in bold and the second-best results are underlined.
@@
3.3 Inference-Time Dynamic Controllability
Although the above pipeline generates natural and expressive motion, the outputs are largely driven by audio, offering limited global control over emotional intensity and spatial dynamics. To address this, we introduce two plug-in controllability modules: global emotion scaling and center motion control, enabling flexible style modulation at inference time.
Global Emotion Scalability. Benefiting from the disentanglement of shape and expression parameters in the FLAME model, we construct expression templates $\{\bar{\boldsymbol{\psi}}^{e}\}_{e=1}^{7}$ from the MEAD dataset, covering seven emotional styles: Angry, Contempt, Disgust, Fear, Happy, Sad, and Surprise. Each template is associated with a template weight $\alpha\in\{1,1.2,1.4,1.6,1.8,2.0\}$ that controls its intrinsic emotional intensity. During inference, we adjust the global emotional tone by interpolating between the reference expression $\boldsymbol{\psi}_{ref}$ and the scaled emotion template:
This provides seven types of global emotion control, each with six levels, while preserving the audio-driven, nuanced expression dynamics. (Details in Appendix  E.1.)
Head-Pose Motion Control. While the predicted mesh sequences exhibit natural but subtle head sway, they lack explicit controllability over head-pose motion. To enhance head presence while preserving realism, we introduce a head-pose control module that provides high-level, interpretable control over head motion. Given the driving audio or a text prompt describing the desired presentation style, a language model generates a simple head-pose trajectory in the form of smooth and interpretable control functions (e.g., gentle sways, rhythmic arcs, or gradual rotations). Rather than replacing the original motion, the generated trajectory is superimposed onto the model-predicted natural head dynamics, yielding controllable yet realistic head-pose motion. This design enables flexible stylistic variation—such as calm, energetic, or stage-presenting delivery—while preserving identity consistency and expression dynamics. Prompt details are provided in Appendix H.
4 Experiments
4.1 Implementation Details
Baselines. We compare our 3DXTalker with seven competitive baselines, including (1) FaceFormer [17]; (2) CodeTalker [50]; (3) SelfTalk [34]; (4) DiffPoseTalk [42]; (5) EMOTE [15]; (6) FaceDiffuser [41]; (7) DEEPTalk [25].
Datasets. We train models on six diverse talking-head datasets, including lab-scanned (GRID [11], RAVDESS [29], MEAD [46]) and in-the-wild sources (CelebV-HQ [53], HDTF [52], and VoxCeleb2 [10]), covering over 11.7k cleaned audio-video pairs with an average of over 15 seconds. We evaluate models on 200 video cases, each with a fixed duration of 10 seconds. More details are in Appendix  B.1.
Setup. The Transformer backbone uses 6 blocks with a hidden size of 768, where each prediction branch has 2 blocks. 3DXTalker is trained for 100 epochs with batch size of 128 and frame size of 250 on two NVIDIA H100 GPUs using AdamW (learning rate $1\times 10^{-4}$, weight decay 0.01) with OneCycleLR scheduler. Flow-matching steps are set to 512 for training and 32 for inference. Implementation details are provided in Appendix  B.2.
Metrics. We evaluate 3DXTalker over 9 metrics across four dimensions. Identity preservation is measured by MVE [16], and CSIM [22], where MVE captures the 3D geometric error between predicted and ground-truth meshes, and CSIM computes frame-wise embedding similarity to ensure identity consistency in 2D appearance. Lip synchronization is assessed using LVE [17], LSEC and LSED [36]; LVE measures 3D vertex-level lip alignment, while LSEC and LSED quantify audio–2D visual sync quality through confidence and embedding-distance mismatch. Facial expression quality is evaluated using UFVE [41], UFDD [50], and Emo-FID [25]; UFVE measures 3D upper-face geometry errors, UFDD examines 3D temporal smoothness of expression dynamics, and Emo-FID reflects 2D facial expression similarities. Head-pose motion is measured using beat alignment (BA) score [42, 40]. Besides, we evaluate perceptual quality via a subjective User Study, reporting the mean rank (MR) based on participant preferences. Finally, FPS is recorded during inference for efficiency evaluation. We compute 3D metrics on meshes and 2D metrics on rendered videos. Details are in Appendix  B.3.
@@ 4.2 Main Results
Quantitative Evaluation. We quantitatively compare our method with seven competitive baselines across identity, lip sync, emotional expression, head pose, and efficiency, listed in Table 1. Our method performs well across the 3D-related metrics (MVE, LVE, UFVE, UFDD), suggesting accurate geometry reconstruction and stable temporal motion compared to the evaluated baselines. For 2D perceptual metrics, 3DXTalker attains competitive results on CSIM and Emo-FID, indicating consistent identity preservation and credible affective quality. The Lip-sync performance is strong overall, while LSEC and LSED remain challenging. We note that most 3D-based methods exhibit similarly limited improvements on these phoneme-sensitive metrics, except for FaceDiffuser. The BA score demonstrates that 3DXTalker can generate head-pose motions that are rhythmically synchronized with the input audio. User study shows that 3DXTalker achieves the best Mean Rank of 4.22, reflecting a human preference for its overall perceptual quality and naturalness compared to the baselines. Finally, 3DXTalker achieves reasonable inference at 69.497 FPS. These results collectively validate that 3DXTalker has strong potential for expressive generation, achieving a balanced trade-off between performance and efficiency.
Figure 3: Qualitative comparisons over selected typical baselines. (a) shows the consistency between generated meshes and the reference image. (b) shows better mouth aperture alignment. (c) shows finer emotional expressiveness. (d) shows predicted natural head pose and camera movements. Full baseline comparisons are provided in Appendix  C. Other emotion comparisons are offered in Appendix  E.2
Table 2: Ablation results of 3DXTalker. “AbsLatent” uses absolute latent instead of differential ones in Eq. (2). “w/o $\mathbf{A}_{\text{emo}}$" and “w/o $\mathbf{A}_{amp}$" remove emotional and amplitude embeddings, respectively. The best results in bold and second-best underlined.
@@
Figure 4: Visualizations of ablation results from Table 2. (a) is conducted on the same audio. (b) extracts each emotion from corresponding videos at the same frame. Details in Appendix  D.
Qualitative Evaluation. Figure 3 provides qualitative comparisons against several representative baselines (Full comparisons are presented in the Appendix  C). As shown in Figure 3 (a), 3DXTalker effectively preserves identity while observing shapes and details, producing meshes that closely match the reference image. In Figure 3 (b), our method achieves clearer mouth aperture alignment across diverse word syllables, demonstrating a tighter correspondence between mouth movements and speech. Figure 3 (c) highlights the model’s ability to generate finer emotional expressions, capturing subtle variations in facial dynamics as the speech evolves. Finally, Figure 3 (d) illustrates natural head-pose behaviors along with different camera movements generated by our 3DXTalker. Overall, these results show that 3DXTalker can synthesize diverse, dynamic, and expressive 3D talking videos.
@@ 4.3 Ablation Study
We conduct ablation experiments to evaluate the contribution of key components in 3DXTalker, with quantitative results listed in Table 2 and visual comparisons shown in Figure 4. Using absolute latent leads to the largest 3D errors and lower CSIM identity scores, indicating that our differential latent design in the dataset-curation pipeline helps separate identity from motion and is important to stable motion generation. Removing $\mathbf{A}_{\text{amp}}$ increases CSIM as it reduces mouth-motion variation, which artificially enhances frame similarity (Figure 4 (a)). It also leads to higher 3D errors, confirming that amplitude cues provide essential constraints on realistic mouth aperture. Likewise, removing $\mathbf{A}_{\text{emo}}$ degrades 3D performance across all geometric metrics and disrupts subtle emotional expression (Figure 4 (b)), underscoring the importance of frame-wise emotion features. These findings validate our incorporation of frame-wise amplitude and emotion cues, forming audio-rich representations that strengthen speech-driven facial motions.
@@
Table 3: Emotion-wise cosine similarity between generated and ground-truth FLAME expression parameters. $E_{C}$" denotes emotion control. Ang. = Angry, Con. = Contempt, Dis. = Disgust, Fea. = Fear, Hap. = Happy, Sad = Sad, Sur. = Surprise. Mesh visualization and analyses are provided in Appendix  E.

### v3

@@ 3.2 Integrated Flow-matching Framework
This branch is responsible for predicting the shape $\{\boldsymbol{\beta}\}$ and detail $\{\boldsymbol{\delta}\}$ parameters that define the identity of the generated talking head. With the global latent representation $\mathbf{H}_{t}$ extracted, the identity head employs lightweight self-attention layers $\mathcal{D}_{id}$ to further form identity-related features and subsequently passes them through an MLP layer to produce the velocity field of shape and detail ($\hat{\boldsymbol{v}}^{\beta}_{t}\in\mathbb{R}^{N\times 100}$ and $\hat{\boldsymbol{v}}^{\delta}\in\mathbb{R}^{N\times 128}$), defined as:
@@
To achieve precise and responsive control of jaw motion and head rotation, we extract frame-wise amplitude features $\mathbf{A}_{amp}$ from the driving audio. Specifically, we first compute the amplitude envelope of the waveform via the Hilbert transform and then apply frame-level window averaging to obtain a temporally aligned amplitude sequence that reflects speech intensity and rhythmic variations. These amplitude cues are injected into the pose branch via a cross-attention module $\mathcal{D}_{pose}$, where $\mathbf{H}_{t}$ provides motion context and $\mathbf{A}_{amp}$ provides audio-driven modulation. The resulting features are decoded by an MLP to produce the velocity fields for jaw pose $\hat{\boldsymbol{v}}^{\theta^{j}}_{t}\in\mathbb{R}^{N\times 3}$ and global head rotation $\hat{\boldsymbol{v}}^{\theta^{g}}_{t}\in\mathbb{R}^{N\times 3}$:
@@
To enable emotionally coherent facial expressions, we extract frame-wise emotion embeddings $\mathbf{A}_{emo}$ from the driving audio using emotion2vec [32]. These embeddings capture subtle affective cues (e.g., happiness, sadness, anger) embedded in speech and are temporally aligned with the latent representation. The expression head injects $\mathbf{A}_{emo}$ into the motion representation via a cross-attention module $\mathcal{D}_{exp}$, followed by an MLP that predicts the expression velocity field $\hat{\boldsymbol{v}}^{\psi}_{t}\in\mathbb{R}^{N\times 50}$, represented as:
@@
Flow-matching Training and Inference. To improve optimization stability in training, we initialize the latent trajectory using a temporally smoothed noise sequence $\boldsymbol{\varepsilon}_{0}\in\mathbb{R}^{T\times D}$. A sparse set of anchor latents is first sampled and then linearly interpolated to form a smooth trajectory. The anchor locations are adaptively determined by extrema of the speech amplitude envelope rather than fixed temporal intervals, introducing a weak articulation-related prior that improves temporal coherence and stabilizes mouth-aperture dynamics. Then, we model the continuous evolution of FLAME parameters as a straight flow between the initial latent $\boldsymbol{\varepsilon}_{0}$ and the target displacement $\mathbf{X}_{\Delta}=\mathbf{X}-\mathbf{X}_{ref}$. Given a randomly sampled $t\sim\mathcal{U}(0,1)$, the target velocity is defined as the linear interpolant, and the flow matching objective supervises the model-predicted velocity $\hat{\boldsymbol{v}}(t)$ as follows:
This formulation directly aligns the model with the continuous flow that transports the latent state toward the target parameter space, enabling stable optimization, temporally smooth synthesis, and coherent evolution across different facial attributes over time.
In inference, the velocity fields from all heads are concatenated ($\boldsymbol{\hat{v}}_{t}=\text{Concat}(\,\hat{\boldsymbol{v}}^{\beta}_{t},\;\hat{\boldsymbol{v}}^{\psi}_{t},\;\hat{\boldsymbol{v}}^{\theta}_{t},\;\hat{\boldsymbol{v}}^{\delta}_{t}\,)$) to form a unified displacement field $\boldsymbol{\varepsilon}{t}$, which is updated iteratively for $T{\text{inf}}$ steps from the initial latent state. The final FLAME parameters are then recovered by adding the predicted displacement to the reference representation:
@@
Overall, the proposed audio-rich flow-matching framework provides a unified approach to model identity and speech-driven facial dynamics in the disentangled FLAME parameter space.
3.3 Decoupled Plug-in Semantic Control
Built upon the proposed dataset and framework, our system already enables expressive and identity-consistent 3D talking avatar generation. However, in real-world scenarios, avatars often require additional global controllability to adapt their delivery styles to different contexts, such as varying emotional intensity or modulating head-pose motion patterns. Although specific control factors can be improved with more targeted data and annotations, broader global stylistic demands are inherently diverse and difficult to enumerate, making it impractical to absorb all such controllability into flow-matching training. Instead, benefiting from the disentangled flow-matching framework, we adopt a decoupled plug-in paradigm that introduces emotion and head-pose control at inference time while preserving learned facial dynamics, and is readily extensible to broader global semantic controls beyond these two cases.
Benefiting from the disentanglement of shape and expression parameters in FLAME, we construct a set of semantic expression templates $\{\bar{\boldsymbol{\psi}}^{e}\}_{e=1}^{7}$ from the MEAD dataset, covering seven emotional styles: Angry, Contempt, Disgust, Fear, Happy, Sad, and Surprise. Each template is associated with an intensity factor $\alpha\in\{1,1.2,1.4,1.6,1.8,2.0\}$ to control the strength of its intrinsic emotional semantics. During inference, we modulate the global emotional semantics by interpolating between the reference expression $\boldsymbol{\psi}_{ref}$ and the scaled emotion template:
This design provides seven emotion categories, each with six controllable intensity levels, while preserving audio-driven expression dynamics. More details are in Appendix D.1.
Complementary to emotion semantic control, we further introduce semantic control over head-pose dynamics. Given the driving audio or a text prompt describing the desired presentation style, a language model generates a stable semantic head-pose trajectory, such as gentle sways, rhythmic arcs, or gradual rotations. This trajectory is superimposed onto the model-predicted natural head motion rather than replacing it, which preserves realism while maintaining stable temporal transitions. In this way, the model supports semantically meaningful stylistic variation, such as calm or energetic delivery, while preserving motion coherence and natural temporal continuity. Details are in Appendix E.1.
Figure 4: Qualitative comparisons over selected typical baselines. (a) shows the consistency between generated meshes and the reference image. (b) shows better mouth aperture alignment. (c) shows finer emotional expressiveness. (d) shows predicted natural head pose and camera movements. Full baseline comparisons are provided in Appendix C, and more emotion comparisons are offered in Appendix D.2. We further present downstream applications of 3DXTalker in Appendix G.
4 Experiments
4.1 Implementation Details
Baselines. We compare 3DXTalker with seven representative baselines spanning diverse paradigms for audio-driven 3D talking-avatar generation. The evaluated baselines are: (1) FaceFormer [16]; (2) CodeTalker [54]; (3) SelfTalk [36]; (4) DiffPoseTalk [45]; (5) EMOTE [14]; (6) FaceDiffuser [44]; and (7) DEEPTalk [27].
Datasets. We train models on six diverse talking-head datasets, covering lab-scanned (GRID [11], RAVDESS [31], MEAD [49]) and in-the-wild sources (CelebV-HQ [57], HDTF [56], and VoxCeleb2 [10]), containing over 11.7k cleaned audio-video pairs with an average of over 15 seconds. We evaluate models on 200 video cases, each with a fixed duration of 10 seconds. More details are in Appendix B.1.
Setup. We use EMOCA with EMOCA_v2_lr_mse_20 model. We use wavlm-base-plus and emotion2vec_plus_base to extract global audio embeddings and frame-wise emotion features from the input audio, respectively. Our integrated audio-rich backbone consists of 6 diffusion transformer blocks with a hidden size of 768, where each prediction head is implemented with 2 blocks. We train 3DXTalker for 100 epochs with a batch size of 128 and a frame length of 250 on two NVIDIA H100 GPUs. Optimization is performed with AdamW (learning rate $1\times 10^{-4}$, weight decay 0.01) with a OneCycleLR scheduler. Flow-matching steps are set to 512 for training and 32 for inference. Finally, following prior works [37, 45], we apply a Savitzky–Golay filter to the predicted sequences, which further improves motion smoothness. More details are in Appendix B.2.
Metrics. We evaluate 3DXTalker using 9 metrics across four dimensions. Identity preservation is measured by MVE [15] and CSIM [22], where MVE reflects 3D geometric error and CSIM measures 2D identity consistency. Lip synchronization is assessed using LVE [16], LSEC and LSED [38], where LVE measures 3D vertex-level lip alignment, while LSEC and LSED quantify audio–2D visual sync quality through confidence and embedding-distance mismatch. Facial expression quality is evaluated using UFVE [44], UFDD [54], and Emo-FID [27]; UFVE measures 3D upper-face geometry errors, UFDD examines 3D temporal smoothness of expression dynamics, and Emo-FID reflects 2D facial expression similarities. Head-pose motion is measured using beat alignment (BA) score [45, 43]. Besides, we evaluate perceptual quality via a subjective User Study, reporting the mean rank (MR) based on participant preferences. Finally, FPS is recorded during inference for efficiency evaluation. We compute 3D metrics on meshes and 2D metrics on rendered videos. Details are provided in Appendix B.3.
@@ 4.1 Implementation Details
Table 1: Quantitative evaluation with baselines over 200 videos. “ID Gen." denotes whether the model supports identity-referred generalization. Our method achieves better performance on 3D-related metrics, demonstrating improved geometric accuracy and temporal coherence, while it performs well on most 2D perceptual metrics. The BA score assesses the rhythmic alignment of head motions with the driving audio. The MR is computed based on voting results from 74 voting responses. Throughput was measured on an NVIDIA RTX 5090 GPU. The best results are highlighted in bold and the second-best results are underlined.
@@ 4.2 Main Results
Qualitative Evaluation. Figure 4 provides qualitative comparisons against several representative baselines (Full comparisons are presented in Appendix C). As shown in Figure 4 (a), 3DXTalker effectively preserves identity while observing shapes and details, producing meshes that closely match the reference image. In Figure 4 (b), our method achieves clearer mouth aperture alignment across diverse word syllables, demonstrating a tighter correspondence between mouth movements and speech. Figure 4 (c) highlights the model’s ability to generate finer emotional expressions, capturing subtle variations in facial dynamics as the speech evolves. Finally, Figure 4 (d) illustrates natural head-pose behaviors along with different camera movements generated by our 3DXTalker. Overall, these results show that 3DXTalker can synthesize diverse, dynamic, and expressive 3D talking videos.
Table 2: Ablation results of 3DXTalker. “AbsLatent” uses absolute latent instead of differential ones in Eq. (2). “w/o $\mathbf{A}_{\text{emo}}$" and “w/o $\mathbf{A}_{amp}$" remove emotional and amplitude embeddings. The best results in bold and second-best underlined.
@@
Figure 5: Visualizations of ablation results from Table 2. (a) is conducted on the same audio. (b) extracts each emotion from corresponding videos at the same frame. More cases can be found in Appendix F.
Quantitative Evaluation. We quantitatively compare our method with seven competitive baselines across identity, lip sync, emotional expression, head pose, and efficiency, as listed in Table 1. Our method performs well across the 3D-related metrics (MVE, LVE, UFVE, UFDD), suggesting accurate geometry reconstruction and stable temporal motion compared to the evaluated baselines. For 2D perceptual metrics, 3DXTalker attains competitive results on CSIM and Emo-FID, indicating consistent identity preservation and credible affective quality. The Lip-sync performance is strong overall, while LSEC and LSED remain challenging. We note that most 3D-based methods exhibit similarly limited improvements on these phoneme-sensitive metrics, except for FaceDiffuser. The BA score demonstrates that 3DXTalker can generate head-pose motions that are rhythmically synchronized with the input audio. User study shows that 3DXTalker achieves the best Mean Rank of 4.22, reflecting a human preference for its overall perceptual quality and naturalness compared to the baselines. Finally, 3DXTalker achieves reasonable inference at 69.497 FPS. These results collectively validate that 3DXTalker has strong potential for expressive generation, achieving a balanced trade-off between performance and efficiency.
@@ 4.3 Ablation Study
We conduct ablation experiments to evaluate the contribution of key components in 3DXTalker, with quantitative results listed in Table 2 and visual comparisons shown in Figure 5. Using absolute latent leads to the largest 3D errors and lower CSIM identity scores, indicating that our differential latent design in the dataset-curation pipeline helps separate identity from motion and is important for stable motion generation. Removing $\mathbf{A}_{\text{amp}}$ increases CSIM since it reduces mouth-motion variation, which artificially enhances frame similarity (Figure 5 (a)). It also leads to higher 3D errors, confirming that amplitude cues provide essential constraints on realistic mouth aperture. Likewise, removing $\mathbf{A}_{\text{emo}}$ degrades 3D performance across all geometric metrics and disrupts subtle emotional expression (Figure 5 (b)), underscoring the importance of frame-wise emotion features. These findings validate our incorporation of frame-wise amplitude and emotion cues, forming audio-rich representations that strengthen speech-driven facial motions.
@@
Table 3: Emotion-wise cosine similarity between generated and ground-truth FLAME expression parameters. $E_{C}$" denotes emotion control. Ang. = Angry, Con. = Contempt, Dis. = Disgust, Fea. = Fear, Hap. = Happy, Sad = Sad, Sur. = Surprise. Mesh visualizations and analyses are offered in Appendix D.
@@
Figure 8: t-SNE visualization of our predicted expression. Partial overlaps between angry–disgust and surprise–fear correspond to their naturally similar facial patterns.
4.4 Analysis
Emotion Analysis. To further verify the model’s emotion controllability, we compute cosine similarity across seven emotions in the FLAME expression space, as listed in Table 3. 3DXTalker consistently achieves higher similarity scores than DEEPTalk, demonstrating superior emotion controllability and coherence. Figure 7 shows temporal emotion controllability. In Figure 7 (a) Without explicit emotion conditioning control, the talking head without emotion condition stays near a neutral state with limited affective variation. In contrast, Figure 7 (b) shows that our hybrid transition control enables smooth, flicker-free transitions across multiple emotion categories while preserving identity-consistent geometry, indicating that the control signal modulates expression in a structured and continuous manner. Moreover, we further present a t-SNE visualization of our expression predictions in Figure 8. Partial overlaps between angry–disgust and surprise–fear are observed due to their similar facial activation patterns, while other clusters remain well separated, indicating that our model captures meaningful structure in the FLAME expression space.
Pose Analysis. We further analyze both jaw and head pose predictions. For jaw pose (Figure 9), the predicted jaw-opening pose ($\boldsymbol{\theta}_{0}^{j}$) trajectories with two individual seeds exhibit temporal consistency with the ground truth and closely follow the audio-amplitude, demonstrating that amplitude conditioning effectively contributes to mouth aperture. Regarding head pose, as shown in Figure 7, 3DXTalker maintains high stability and continuity across both supported modes. The results demonstrate that the model achieves both realistic pose generation and controllable dynamics without introducing abrupt jitter or pose discontinuities.